OpenStudy (anonymous):

The volume V of a given mass of gas varies directly as the temperature T and inversely as the pressure P. If V=180 in.^3 when T=150 degrees and P=10lb/in^2, what is the volume when T=400 degrees and P=10lb/ in.^2?

OpenStudy (anonymous):

So this is a chemistry question that is easily solved by the ideal gas law. \[PV/T=PV/T\] Luckily, all the measurements are in the same units, so we must continue to obey them. \[10*180/150=10*V/400\] What can be noted is that the pressures are the same, so we could have stuck with the simpler Charles' Law, \[V/T=V/T\] V=480 in^3